a multicab as a velocity of 18 point 52 km per hour going 30 ยฐ west of north what are its horizontal and vertical components
1
Expert's answer
2021-10-13T09:06:45-0400

Calculate the horizontal (westward) component:


"v_W=v\sin\theta=18.52·\sin30°=9.26\text{ km/h}."

The vertical (northward) component:


"v_N=v\cos\theta=18.52·\cos30°=16.04\text{ km/h}."